can anyone help
Find the equation, in standard form, with all integer coefficients, of the line perpendicular to x – 2y = 24 and passing through (1, 3).

Answer by Alan3354(69443)

can anyone help
Find the equation, in standard form, with all integer coefficients, of the line perpendicular to x – 2y = 24 and passing through (1, 3).
--------------------
Perpendicular means the slope is the negative inverse of the slope of the given line.
1st, find the slope of the given line by putting it into slope-intercept form, y = mx+b
------------
x – 2y = 24
-2y = -x + 24
y = (1/2)x - 12
The slope, m, is 1/2. So the slope of the line perpendicular will have a slope of -2.
--------
Use the slope-intercept again:
y-y1 = m*(x-x1) *** (x1,y1) is the point (1,3)
y-3 = 2*(x-1)
y-3 = 2x-2
2x - y = -1 (standard form)